Bonds payable, 10% (issued at face amount) $2,500,000 Preferred $2 stock, $20 par 2,500,000 Common stock, $25 par 2,500,000 Income tax is estimated at 40% of income. Determine the earnings per share on common stock, assuming that the income before bond interest and income tax is (a) $1,125,000, (b) $1,375,000, and (c) $1,625,000. Enter answers in dollars and cents, rounding to two decimal places. a. Earnings per share on common stock $ b. Earnings per share on common stock $ c. Earnings per share on common stock $
Answer #1

Answer:

$2.75 Earnings per share on common stock Earnings per share on common stock $4.25 $5.75 Earnings per share on common stock

Working:

Interest expense = 2500000 * 10% = $250,000

Preference dividend = $2/20 * 2500000 = $250,000

Number of common stock outstanding = 2500000 / 25 = 100000

Assumptions: (a) $1,125,000 $250,000 $875,000 $350,000 (b) $1,375,000 $250,000 (c) $1,625,000 Income before interest and Taxe